Find an angle such that 3 times the complement of the angle is 50 degrees greater than the supplement of the angle.
Answer by nerdybill(7384)

Find an angle such that 3 times the complement of the angle is 50 degrees greater than the supplement of the angle.
.
Two things you need to know:
If two angles are supplementary, the sum of the two angles = 180.
If two angles are complementary, the sum of the two angles = 90.
.
Let x = the angle
then
90-x = complement of angle
180-x = supplement of angle
.
3(90-x) = (180-x) + 50
270-3x = 180-x+50
270-3x = 230-x
270 = 230+2x
40 = 2x
20 degrees = x